# cp3-bench: A tool for benchmarking symbolic regression algorithms tested with cosmology

### v1

@@ 3.2 Benchmarking results: Cosmological datasets
In this subsection we present the results from benchmarking the algorithms on our 19 cosmological datasets.
@@
Before discussing the results, we provide some comments regarding the presented MSE values and their reported equations: The algorithms considered are created by other authors and can come with unique features, and thus they behave in different ways. Therefore, some algorithms allow you to define thresholds for dropping certain residue terms, and some do these simplification out of the box without user input. Therefore, different algorithms do certain simplifications including rounding numbers to some number of significant digits.
When we compute the MSEs presented in the forthcoming tables, we use the full symbolic expression found by the algorithm. We call this the reported MSE. As explained above, the full symbolic expression is sometimes not identical to the reported expression (the expression printed out by the algorithm). This is simply because the reported expression may, as noted above, have residue terms drops and presented with less significant digits in order to make the output equation more presentable.
@@
As part of the result processing we have extracted the reported equations into mathematica notebooks where we have recomputed the MSE based on the reported equation. We find that due to the mentioned simplifications, the reported MSE does not always match the MSE computed from the reported equation. We emphasize that the MSE values reported in this paper are those coming from the full symbolic expression identified by the given algorithm. This is also the MSE which is provided by cp3-bench.
@@
For most algorithms we do not find significant deviations between the MSE computed using the full expression and the reported expression. However, there are some exceptions worth mentioning. We first note that GPZGD consistently showed a significantly lower reported MSE compared to the MSE of the reported expression. This was also the case for FFX for the more complicated equations corresponding to the datasets $C4$, $C6$, $C7$, $C12$, $C13$, $C16$, and $C17$. We attribute this to the fact that FFX only outputs 3 significant digits in the output equation, but the underlying equation identified by the algorithm can have more significant digits. The resulting differences between the reported MSE and the MSE obtained from the reported expression deviate more for more complicated and long equations. We further note that Genetic Engine and AI-Feynman do not protect against producing imaginary numbers. The results obtained from Genetic Engine on dataset $C19$ and from AI Feynman on dataset $C6$ for instance contained imaginary parts. In these cases, only the real value is reported by the algorithm. Such situations can also lead to differences in the reported MSE and the MSE of the reported algorithm. We also note that the reported equation of Genetic Engine would yield division by zero errors for one or more parameters considered for the datasets $C6$, $C8$, $C15$, $C16$ and $C17$. This division by zero would presumably not appear in the exact expressions obtained before rounding etc. In general, the reported MSE and the MSE based on the reported equation were for QLattice very close to each other. The exception is for $C11$ where the reported equation produces an MSE that is orders of magnitude worse than the reported MSE based on the exact expression. Therefore, we have put this value in red in Tables 4 and LABEL:table:best_results.
@@
Cosmological dataset results Algorithm Dataset AI-Feynman DSO DSR FFX Genetic Engine GPZGD ITEA PySR QLattice uDSR C1 $10^{-8}$ $10^{-5}$ $10^{-5}$ $10^{-6}$ $10^{-6}$ $10^{-13}$ $10^{-14}$ $10^{-9}$ $10^{-10}$ C2 $10^{-14}$ $10^{-14}$ $10^{-14}$ $10^{-14}$ $10^{-14}$ $10^{-14}$ $10^{-14}$ $10^{-14}$ $10^{-14}$ $10^{-14}$ C3 $10^{-6}$ $10^{-4}$ $10^{-4}$ $10^{-6}$ $10^{-5}$ $10^{-6}$ $10^{-9}$ $10^{-6}$ $10^{-9}$ C4 $10^{-5}$ $10^{-6}$ $10^{-5}$ $10^{-5}$ $10^{-6}$ $10^{-5}$ $10^{-8}$ $10^{-6}$ $10^{-7}$ C5 $10^{-4}$ $10^{-4}$ $10^{-5}$($10^{-31}$) C6 $10^{-3}$ $10^{-4}$ $10^{-4}$ $10^{-6}$ $10^{-5}$ $10^{-7}$ $10^{-7}$ $10^{-5}$ $10^{-8}$ C7 $10^{-6}$ $10^{-4}$ $10^{-5}$ $10^{-4}$ $10^{-6}$ $10^{-10}$ $10^{-13}$ $10^{-7}$ $10^{-11}$ $10^{-13}$ C8 $10^{-4}$ $10^{-3}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-7}$ $10^{-10}$ $10^{-6}$ $10^{-9}$ $10^{-11}$ C9 $10^{-3}$ $10^{-3}$ $10^{-3}$ $10^{-4}$ $10^{-4}$ $10^{-5}$ $10^{-6}$ $10^{-3}$ $10^{-6}$ $10^{-7}$ C10 $10^{-3}$ $10^{-2}$ $10^{-2}$ $10^{-3}$ $10^{-3}$ $10^{-3}$ $10^{-5}$ $10^{-2}$ $10^{-5}$ $10^{-6}$ C11 $10^{-3}$ $10^{-3}$ $10^{-3}$ $10^{-4}$ $10^{-3}$ $10^{-4}$ $10^{-7}$ $10^{-3}$ $10^{-6}$ $10^{-3}$ C12 $10^{-4}$ $10^{-2}$ $10^{-3}$ $10^{-3}$ $10^{-2}$ $10^{-1}$ $10^{-6}$ $10^{-3}$ $10^{-6}$ $10^{-2}$ C13 $10^{-3}$ $10^{-2}$ $10^{-3}$ $10^{-4}$ $10^{-4}$ $10^{-3}$ $10^{-5}$ $10^{-4}$ $10^{-6}$ $10^{-2}$ C14 $10^{-4}$ $10^{-3}$ $10^{-3}$ $10^{-4}$ $10^{-4}$ $10^{-5}$ $10^{-2}$ $10^{-3}$ $10^{-6}$ $10^{-3}$ C15 10 10 10 10 10 10 10 10 10 10 C16 1 $10^{-1}$ $10^{-1}$ $10^{-1}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ C17 $10^{-1}$ 1 $10^{-1}$ $10^{-1}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ 1 C18 $1$ $10^{-1}$ $10^{-1}$ $10^{-2}$ $10^{-2}$ $10^{-3}$ $10^{-3}$ $10^{-2}$ $10^{-3}$ 1 C19 $10^{-3}$ $10^{-4}$ $10^{-3}$ $10^{-3}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-2}$
Table 4: Result of evaluating the cosmological datasets across all algorithms.

### v2

@@ 3.2 Benchmarking results: Cosmological datasets
In this subsection we present the results from benchmarking the algorithms on our 28 cosmological datasets.
@@
Before discussing the results, we provide some comments regarding the presented MSE values and their reported equations: The algorithms considered are created by other authors and can come with unique features, and thus the algorithms behave differently from each other. For instance, some algorithms allow you to define thresholds for dropping certain residue terms, and some algorithms even do these simplifications out-of-the-box without user input. This means that the symbolic expression outputs can deviate from the “un-simplified” expressions that the algorithms use for computing MSEs.
@@
When we compute the MSEs presented in the forthcoming tables, we use the full symbolic expression found by the algorithm, before the algorithm simplifies the expression by rounding to a certain number of digits etc. We will refer to this at the reported MSE. The full symbolic expression is not always identical to the reported expression (the expression printed out by the algorithm). This is because the reported expression may, as noted above, have residue terms dropped and presented with less significant digits in order to make the output equation more presentable. As part of the result-processing we have extracted the reported equations into Mathematica notebooks where we have recomputed the MSE based on the reported equation. These notebooks can be found in the Things-to-bench repository. We find that due to the mentioned simplifications, the reported MSE does not always match the MSE computed from the reported equation. We emphasize that the MSE values reported in this paper are those coming from the full symbolic expression identified by the given algorithm. This is also the MSE which is provided by cp3-bench.
@@
For most algorithms we do not find significant deviations between the MSE computed using the full expression and the reported expression. However, there are some exceptions worth mentioning. We first note that GPZGD consistently showed a significantly lower reported MSE compared to the MSE of the reported expression. This was also the case for FFX for the more complicated equations corresponding to the datasets $C1d$, $C2b$, $C3a$, $C3f$, $C3g$, $C4a$, $C4c$, $C5c$, $C5d$, $C5e$, $C5f$, $C6a$ and $C6b$. We attribute this to the fact that FFX only outputs 3 significant digits in the output equation, but the underlying equation identified by the algorithm can have more significant digits. The resulting differences between the reported MSE and the MSE obtained from the reported expression thus deviate more for more complicated and long expressions. We further note that Genetic Engine and AI-Feynman do not protect against producing imaginary numbers. The results obtained from Genetic Engine on dataset $C4e$ and from AI Feynman on dataset $C2b$ and $C6c$ for instance contained imaginary parts. In these cases, only the real value is reported by the algorithm. Such situations can also lead to differences in the reported MSE and the MSE of the reported algorithm. We also note that the reported equation of Genetic Engine would yield division by zero errors for one or more parameters considered for the datasets $C2b$, $C3b$, $C4a$, $C4b$, $C4c$ and $C6b$. This division by zero would presumably not appear in the exact expressions obtained before rounding etc. In general, the reported MSE and the MSE based on the reported equation were for ITEA very similar. The only exception is for $C3e$ where the reported equation produces an MSE that is orders of magnitude worse than the reported MSE based on the exact expression. Therefore, we have put this value in red in Tables 4 and LABEL:table:best_results.
@@
Cosmological dataset results Algorithm Dataset AI-Feynman DSO DSR FFX GE GPG GPZGD ITEA Operon PySR QLattice uDSR C1a $10^{-8}$ $10^{-5}$ $10^{-5}$ $10^{-6}$ $10^{-6}$ $10^{-2}$ $10^{-13}$ $10^{-14}$ $10^{-5}$ $10^{-9}$ $10^{-10}$ C1b $10^{-14}$ $10^{-14}$ $10^{-14}$ $10^{-14}$ $10^{-14}$ $10^{-3}$ $10^{-14}$ $10^{-14}$ $10^{-5}$ $10^{-14}$ $10^{-14}$ $10^{-14}$ C1c $10^{-6}$ $10^{-4}$ $10^{-4}$ $10^{-6}$ $10^{-5}$ $10^{-8}$ $10^{-6}$ $10^{-9}$ $10^{-4}$ $10^{-6}$ $10^{-9}$ C1d $10^{-5}$ $10^{-6}$ $10^{-5}$ $10^{-5}$ $10^{-6}$ $10^{-6}$ $10^{-5}$ $10^{-8}$ $10^{-4}$ $10^{-6}$ $10^{-7}$ C2a $10^{-4}$ $10^{-4}$ $10^{-5}$($10^{-31}$) $10^{-3}$ C2b $10^{-3}$ $10^{-4}$ $10^{-4}$ $10^{-6}$ $10^{-5}$ $10^{-7}$ $10^{-7}$ $10^{-7}$ $10^{-3}$ $10^{-5}$ $10^{-8}$ C3a $10^{-6}$ $10^{-4}$ $10^{-5}$ $10^{-4}$ $10^{-6}$ $10^{-1}$ $10^{-10}$ $10^{-13}$ $10^{-4}$ $10^{-7}$ $10^{-11}$ $10^{-13}$ C3b $10^{-4}$ $10^{-3}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{0}$ $10^{-7}$ $10^{-10}$ $10^{-4}$ $10^{-6}$ $10^{-9}$ $10^{-11}$ C3c $10^{-3}$ $10^{-3}$ $10^{-3}$ $10^{-4}$ $10^{-4}$ $10^{-5}$ $10^{-5}$ $10^{-6}$ $10^{-3}$ $10^{-3}$ $10^{-6}$ $10^{-7}$ C3d $10^{-3}$ $10^{-2}$ $10^{-2}$ $10^{-3}$ $10^{-3}$ $10^{-5}$ $10^{-3}$ $10^{-5}$ $10^{-3}$ $10^{-2}$ $10^{-5}$ $10^{-6}$ C3e $10^{-3}$ $10^{-3}$ $10^{-3}$ $10^{-4}$ $10^{-3}$ $10^{-5}$ $10^{-4}$ $10^{-7}$ $10^{-3}$ $10^{-3}$ $10^{-6}$ $10^{-3}$ C3f $10^{-4}$ $10^{-2}$ $10^{-3}$ $10^{-3}$ $10^{-2}$ $10^{-5}$ $10^{-1}$ $10^{-6}$ $10^{-3}$ $10^{-3}$ $10^{-6}$ $10^{-2}$ C3g $10^{-3}$ $10^{-2}$ $10^{-3}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-3}$ $10^{-5}$ $10^{-3}$ $10^{-4}$ $10^{-6}$ $10^{-2}$ C3h $10^{-4}$ $10^{-3}$ $10^{-3}$ $10^{-4}$ $10^{-4}$ $10^{-5}$ $10^{-5}$ $10^{-2}$ $10^{-3}$ $10^{-3}$ $10^{-6}$ $10^{-3}$ C4a $10^{-2}$ $10^{-1}$ $10^{-1}$ $10^{-1}$ $10^{-2}$ $10^{2}$ $10^{-1}$ $10^{-2}$ $10^{-2}$ $10^{-1}$ $10^{-2}$ $10^{-2}$ C4b $10^{0}$ $10^{-1}$ $10^{-1}$ $10^{-1}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ C4c $10^{-1}$ $10^{0}$ $10^{-1}$ $10^{-1}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{0}$ C4d $10^{0}$ $10^{-1}$ $10^{-1}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-3}$ $10^{-3}$ $10^{-2}$ $10^{-2}$ $10^{-3}$ $10^{0}$ C4e $10^{-3}$ $10^{-4}$ $10^{-3}$ $10^{-3}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-5}$ $10^{-4}$ $10^{-4}$ $10^{-2}$ C5a (✓) (✓) $10^{-2}$ (✓) (✓) ($10^{-13}$) $10^{-1}$ $10^{-9}$ $10^{-4}$ (✓) ✓($10^{-16}$) $10^{-4}$ C5b (✓) $10^{-4}$ $10^{-2}$ $10^{-4}$ $10^{-4}$ $10^{-1}$ $10^{-4}$ $10^{-4}$ (✓) ✓($10^{-4}$) $10^{-3}$ C5c $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-1}$ $10^{-1}$ $10^{-2}$ $10^{-2}$ (✓) $10^{-2}$ $10^{-2}$ C5d $10^{0}$ $10^{-3}$ $10^{-3}$ $10^{-2}$ $10^{-3}$ $10^{-6}$ $10^{-1}$ $10^{-7}$ $10^{-3}$ $10^{-4}$ $10^{-2}$ C5e $10^{-1}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-7}$ $10^{-4}$ $10^{-7}$ $10^{-5}$ $10^{-8}$ $10^{-3}$ C5f $10^{0}$ $10^{-2}$ $10^{-2}$ $10^{-1}$ $10^{-2}$ $10^{-5}$ $10^{-1}$ $10^{-3}$ $10^{-3}$ $10^{-2}$ $10^{-3}$ $10^{-1}$ C6a $10^{-1}$ $10^{-1}$ $10^{-1}$ $10^{-1}$ $10^{-1}$ $10^{-1}$ $10^{-2}$ $10^{-1}$ $10^{-2}$ $10^{-1}$ $10^{-1}$ $10^{-1}$ C6b $10^{-3}$ $10^{-1}$ $10^{-1}$ $10^{-1}$ $10^{-1}$ $10^{-5}$ $10^{-5}$ $10^{-3}$ $10^{-2}$ $10^{-3}$ $10^{-2}$ $10^{-1}$ C6c $10^{0}$ $10^{0}$ $10^{0}$ $10^{0}$ $10^{0}$ $10^{-1}$ $10^{0}$ $10^{0}$ $10^{0}$ $10^{0}$ $10^{0}$ $10^{0}$
Table 4: Result of evaluating the cosmological datasets across all algorithms. We highlight a single MSE in red because the reported MSE (shown in the table) deviates significantly from the MSE of the reported equation. The green background colors indicate a correct symbolic expression, orange is an almost correct expression, and the different red colors indicate the size of the MSE.